# Expressing and visualizing model uncertainty in Bayesian variable selection using Cartesian credible sets

### v2

@@ 7.1 GWAS example: Systemic Lupus Erythematosus
We consider a GWAS for Systemic Lupus Erythematosus using a case-control study. The presence or absence of the disease is the binary outcome in a logistic regression with the SNPs as explanatory variables. It is assumed that there are a number of causal SNPs which truly cause biological differences between the case and controls. A simple analysis would fit a logistic regression separately for each SNP and declare a SNP as associated with the outcome if it has a p-value smaller than a chosen (typically, small) threshold. However, due to correlation in the genome, we usually see many SNPs declared as associated for each causal SNP. We refer to this group of SNPs as the signal. The goal of the analysis is to determine the number of independent signals and which SNPs are most likely the causal SNPs at each signal.
@@
The data contains 4,036 cases and 6,959 controls, with the cases and controls having European ancestry. In this example we only look at SNPs on Chromosome 3 ($42,430$ SNPs/variables). We apply Bayesian variable selection to the data using a logistic regression model with all SNPs as potential covariates and fit the model using an adaptive MCMC algorithm (Griffin and Wan, 2021). The full specification of the Bayesian model is provided in Appendix C. Figure 3 shows the PIPs for each SNP. There are several clear signals from SNPs with PIPs equal to 1 and others with large PIPs. The 50% Cartesian credible sets are found by first retaining only those SNPs with a PIP above 0.04 (leaving 69 variables). The genomic distances of these 69 variables are given in Appendix C. The median model contains 13 variables which using the indices of the retained SNPs are 5, 10, 37, 45, 47, 48, 50, 51, 52, 53, 54, 58, 63.
@@
Figure 4: GWAS example: Thresholding Method, $M=2$. Black lines indicate the blocks of the partition.
@@
The 50% Cartesian credible sets found using the thresholding method and the agglomerative method are shown in Figures 4 and 5 respectively. The most striking aspect is that the blocks of variables are contiguous which reflects the serial correlation of the SNPs. The results using the thresholding method in Figure 4 show that 11 variables in the median model do not include the possibility of exclusion in the credible set (10, 45, 47, 48, 50, 51, 52, 53, 54, 58, 63). Two of these variables are included as a block: 58 with 55, 56 and 57, and 63 with 64, 65, 66 and 67. There are two blocks of variables which have no variables included in the median model but which have substantial support for inclusion as a block. These are 32, 33, 34, 35 and 36 (genomic distances: 103387440–103433922), which has a posterior probability of 0.61 of including one of the variables (the largest PIP for a variable in the block is 0.23), and 59, 60, 61, and 62 (genomic distance: 159767453–159791628), which has a posterior probability of 0.84 (the largest PIP for variables in the block is 0.28).
Figure 5: GWAS example: Agglomerative Method, $M=2$. Black lines indicate the blocks of the partition.
The results from the agglomerative method in Figure 5 are very similar to the results using the thresholding method but are more conservative with only the extra blocks $\{59,60,61,62\}$ and $\{63,64,65,66,67\}$ included. The method groups 57 with 58. The variables 32, 33, 34, 35 and 36 are not included as a block. The difference is marginal with the block $\{32,33,34,35,36\}$ from the thresholding method having fairly weak support to be included as a block.
7.2 Models with interactions: Concrete data
Schäfer and Chopin (2013) describe several data sets designed to have high-levels of multi-collinearity to test computational methods for sampling posterior distribution from Bayesian variable selection. We concentrate on their Concrete data example, which has 1030 observations and eight variables, which are listed in Appendix D. The reader is referred to their paper for full details of the regression problem. They construct a challenging variable selection problem by including transformation of the recorded covariates and two-way interactions. Table 3 shows the 13 variables created by transforming these original variables.
$X_{1}$ $\log X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $\log X_{4}$ $X_{5}$ $X_{6}$ $\log X_{6}$ $X_{7}$ $\log X_{7}$ $X_{8}$ $\log X_{8}$
1 2 3 4 5 6 7 8 9 10 11 12 13
Table 3: Concrete examples variables
Figure 6: Concrete data: Thresholding Method using prior I, $M=2$. Black lines indicate the blocks of the partition.
We consider how the choice of prior affects the model uncertainty through the lens of 50% Cartesian credible sets. We fit a linear regression which includes the 13 main effects and 78 two-way interactions, and consider two choices of prior. Firstly, we ignore the distinction between main effects and two-way interaction and treat this as a variable selection problem with 91 regressors. We use the hierarchical prior in Appendix B, which we term Prior I. An alternative prior can be constructed which distinguishes between main effects and interaction using the idea of strong heredity (that interactions can only be included in a model if both main effects are included). These types of prior were introduced in Chipman (1996). We use the prior in Appendix B but with the prior probability of 0.5 of including an interaction (conditional on including both main effects) and we give the prior of including a main effect, $\pi$, the prior $\pi\sim\mbox{Be}\left(1,\frac{p-3p_{0}+\sqrt{(p_{0}+p)^{2}+2pp_{0}(p-1)}}{2p_{0}}\right)$ where $p$ is the number of variables and $p_{0}$ is the prior mean number of included variables. The choice of prior for $\pi$ implies that the prior mean number of included variables is $pp_{0}$ as in Prior I. This is called Prior II. Both priors imply the same level of prior sparsity but differ in the structure of prior. With Prior I we used the MCMC set-up for the George & McCulloch example described in Appendix B and for Prior II we used an extension of the ASI algorithm (Griffin et al., 2021) to regression models with two-way interactions described in Appendix E
Figure 7: Concrete data: Thresholding Method using prior II, $M=2$. Black lines indicate the blocks of the partition.
The results of using the thresholding method with Priors I and II are shown in Figures 6 and 7 (the results using the Agglomerative Method are provided in Appendix D). The results with the two priors are strikingly different. Prior I leads to a much larger 50% Cartesian credible set which includes $3.98\times 10^{21}$ models whereas Prior II leads to a set with $4.95\times 10^{13}$. The set using Prior I is 80 000 000 times larger than the set using Prior II. This shows the ability of a well chosen prior to substantially reduce the amount of model uncertainty. Prior II also leads to much clearer results with an inclusion variable equal to 1 for 25 variables compared to 10 variables under Prior II.
@@ 8 Discussion
Cartesian credible sets are a useful way to summarize and visualize the output from Bayesian variable selection. The sets are constructed by finding a partition of the variable and an approximation to the posterior distribution where variables in different subsets are independent. Credible sets are defined as a Cartesian products of sets of sub-model and can be found using a simple heuristic algorithm. Two methods for finding the approximation are proposed: 1) thresholding the correlation matrix and 2) an agglomerative algorithm for minimizing the KL divergence between the posterior and the approximation. In both cases, there is a parameter controlling the level of approximation and this can be chosen using a criterion which trades-off the accuracy of the approximation and overall the size of the credible set. The method can be run quickly as post-processing of an MCMC sample from the posterior distribution on the models. The results can be visualized by plotting the list of sub-models for each subset and coloured according to the posterior probability of the sub-models and the posterior probability of including at least one variable in a subset. This gives a much more complete picture of the posterior distribution than the usually plotted PIPs and helps understanding of the interchangeability of variables which are highly collinear. This allows us to avoid dilution effects from posterior mass being distributed across such variables by considering blocks of variables.
@@ A.1 Proposition 1
Appendix B George and McCulloch (1997) example
@@
The posterior distribution was sampled using a simple Add-Delete-Swap sampler for $\gamma$ with Gibbs steps to update $\pi$ and $g$ and all other parameters integrated out. The samplers were run for 60 000 iterations with the first 10 000 iterations used as a burn-in, and the subsequent 50 000 iterations thinned every fifth value to leave a final sample of 10 000 values.
Appendix C GWAS data
@@
The posterior distribution was sampled using the ASI sampler for logistic regression from Griffin and Wan (2021). Three independent ASI chains were run with each chain containing 25 replicate chains. Each ASI was run for 15 000 iterations with 5 000 iterations used a burn-in and the subsequent 10 000 samples thinned every 10th value. This gave 1 000 samples for each replicate chain, 25 000 samples for each ASI chain, and an overall number of 75 000 posterior samples.
@@
Appendix D Concrete data
The original variables in the Concrete data are shown in Table 5. The results of using the Agglomerative Method with Prior I and Prior II are shown in Figures 8 and 9 respectively. The credible sets are similar to the corresponding ones using the Thresholding Method but are slightly more conservative.
Original variables $X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ $X_{6}$ $X_{7}$ $X_{8}$ Cement Blast Furnace Fly Water Superplasticizer Coarse Fine Age Slag Ash Aggregate Aggregate
Table 5: Concrete examples original variables
Figure 8: Concrete data: Agglomerative Method with prior I, $M=2$. Black lines indicate the blocks of the partition.
Figure 9: Concrete data: Agglomerative Method with prior II, $M=2$. Black lines indicate the blocks of the partition.
Appendix E ASI algorithm for a linear model with two-way interactions
We are interested in a model with main effects and interactions. Let $X$ be a design matrix for the main effects, $Z$ be a design matrix for the interactions and let $b_{i,1}$ and $b_{i,2}$ be the main effects that correspond to the $i$-th interaction. We assume the linear model
$Y=X_{\gamma}+Z_{\kappa}+\epsilon$
where $\epsilon\sim\mbox{N}(0,\sigma^{2}I)$, $\gamma$ and $\kappa$ are the inclusion variables for the main effects and interactions respectively, $X_{\gamma}$ and $Z_{\kappa}$ are the corresponding design matrices for the included main effects and included interactions respectively. The prior inclusion probability are given a weak heredity prior which has the form
$p(\gamma_{j})=\pi_{M}$
$p(\kappa_{i}=1|\gamma_{b_{i,1}},\gamma_{b_{i,2}})=\left\{\begin{array}[]{ll}\pi_{1}&\mbox{if }\gamma_{b_{i,1}}=0,\gamma_{b_{i,2}}=0\\ \pi_{2}&\mbox{if }\gamma_{b_{i,1}}=0,\gamma_{b_{i,2}}=1\mbox{ or }\gamma_{b_{i,1}}=1,\gamma_{b_{i,2}}=0\\ \pi_{3}&\mbox{if }\gamma_{b_{i,1}}=1,\gamma_{b_{i,2}}=1\end{array}\right.$
where $\pi_{1}<\pi_{2}<\pi_{3}$. Strong heredity (i.e. the idea that an interaction can only be included in the variable if both its main effects are included) corresponds to $\pi_{1}=\pi_{2}=0$. The ASI sampler(Griffin et al., 2021) increasingly struggles to mix as $\pi_{1}$ and $\pi_{2}$ become closer to zero since the posterior distribution will show increasing correlation between interactions and the corresponding main effects.
The ASI sampler can be adapted by introducing a two-step proposal which has the form
$q_{main}(\gamma,\gamma^{\prime})=\prod_{i=1}^{p}A_{i}^{(\gamma_{i}-1)\gamma_{i}^{\prime}}(1-A_{i})^{(\gamma_{i}-1)(\gamma_{i}^{\prime}-1)}D_{i}^{\gamma_{i}(1-\gamma_{i}^{\prime})}(1-D_{i})^{\gamma_{i}\gamma_{i}^{\prime}}$
where $A_{i}$ and $D_{i}$ are the calculated using the Rao-Blackweillised estimator described in Griffin et al. (2021). The interactions are proposed (conditional on $\gamma$ and $\gamma^{\prime}$) from
$q_{int}(\kappa,\kappa^{\prime})=\prod_{i=1}^{q}\tilde{A}_{i}^{(\kappa_{i}-1)\kappa_{i}^{\prime}}(1-\tilde{A}_{i})^{(\kappa_{i}-1)(\kappa_{i}^{\prime}-1)}\tilde{D}_{i}^{\kappa_{i}(1-\kappa_{i}^{\prime})}(1-\tilde{D}_{i})^{\kappa_{i}\kappa_{i}^{\prime}}$
where $\tilde{A}_{i}$ and $\tilde{D}_{i}$ by
$\tilde{A}_{i}=\zeta\min\left\{1,\frac{\tilde{\pi}_{i,\gamma^{\prime}_{b_{i,1}}+\gamma^{\prime}_{b_{i,2}}}}{1-\tilde{\pi}_{i,\gamma_{b_{i,1}}+\gamma_{b_{i,2}}}}\right\}$
and
$\tilde{D}_{i}=\zeta\min\left\{1,\frac{1-\tilde{\pi}_{i,\gamma^{\prime}_{b_{i,1}}+\gamma^{\prime}_{b_{i,2}}}}{\tilde{\pi}_{i,\gamma_{b_{i,1}}+\gamma_{b_{i,2}}}}\right\}.$
The parameter $\zeta$ is common to both proposals. The PIPs for the interactions in the proposal $\tilde{\pi}_{i,j}$ are Rao-Blackwellised estimates of the PIP $p(\kappa_{i}=1|\gamma_{b_{i,1}}+\gamma_{b_{i,2}}=j)$. This allows the proposal for the interaction terms to adapt to the included/excluded main effects. This becomes increasingly important if $\pi_{1}$ and $\pi_{2}$ are close to zero. In the limit, in the case of strong heredity, $p(\kappa_{i}=1|\gamma_{b_{i,1}}+\gamma_{b_{i,2}}<2)=0$ and so must be different from $p(\kappa_{i}=1|\gamma_{b_{i,1}}+\gamma_{b_{i,2}}=2)$. These probabilities will also be very different if $\pi_{1}$ and $\pi_{2}$ are close to zero. The proposal is accepted using the usual Metropolis-Hastings ratio.

### v3

@@ 7 Applications
We consider four applications to illustrate the use of Cartesian credible sets in real problems. The first three examples use linear regression models to analyse data from the UCI Machine Learning Repository and the fourth example uses a logistic regression model to analyse data from a genome-wide association study (GWAS) of Systemic Lupus Erythematosus. The prior settings and MCMC schemes used for inference are given in Appendix B.
7.1 Behaviour of traffic in Sao Paulo, Brazil
The data describe traffic levels in Sao Paulo, Brazil from 14th to 18th December 2009.
Figure 5: Sao Paulo traffic example: $50\%$ Cartesian credible set with $M=2$. Black lines indicate the blocks of the partition.
A measure of the slowness of traffic is recorded from 7am to 8pm at half-hour intervals giving 27 measurements for each day. There are also 16 covariates recorded about problems in the Sao Paulo traffic network. We fit a regression model which uses a random walk to model the effect of the hour-of-day, the 16 covariates and an effect for the day of the week. The regression model for the slowness of traffic of the $i$-th observation, $y_{i}$, is
$y_{i}=m(x_{i})+\sum_{i=27}^{46}\beta_{i}Z_{i}+\epsilon_{i}$
where $x_{i}$ is the hour, $Z_{i}$ are the covariates and day-of-the-week dummy variables and $\epsilon_{i}\sim\mbox{N}(0,\sigma^{2})$. We model $m(x)$ using a random walk giving regression coefficients $\beta_{i}=m(i+1)-m(i)$ for $i=1,\dots,26$, for the $i$-th observation.
The 50% Cartesian credible set is shown in Figure 5. There are three clear blocks of variables with blocks PIPs above 0.5 (those with variables coloured in pink). The first two blocks (2, 3, 4) and (20, 21, 22) correspond to increments in the random walk prior for $m(x)$. Figure 6 shows the mean slowness in traffic over time.
Figure 6: Sao Paulo traffic example: mean Slowness in Traffic as a function of time. The grey blocks indicate the times associated with regressors (2, 3, 4) and (20, 21, 22).
The times associated with these two blocks of regressors are indicated by grey shading and, clearly, these are the times with the largest increases in slowness of traffic. The uncertainty in the variable selection for these blocks represents uncertainty about which and how many increments are needed to model these changes in the level of the response. The third block has a high posterior probability of including at least one of the variables (0.97) but there’s substantial uncertainty.
7.2 Yield of soybean
soybean collected data from an experiment to understand the effect of physical characteristics of cultivars of soybean on grain yields. The experiment was run in two seasons with four replicates of 40 cultivars giving a total of 320 observations. We use a linear regression model to estimates the effects of the season and the characteristics.
Variable Description Variable Description
S Season (1 or 2) NGP Number of legumes per plant (unit)
PH plant height (cm) NGL Number of grains per plant (unit)
IFP insertion of the first pod (cm) NS Number of grains per pod (unit)
NLP Number of stems (unit) MHG Thousand seed weight (g)
Table 3: Soybean example: the variables in the regression
A list of the variables are given in Table 3.
Figure 7: Soybean example: $50\%$ CCS with $M=2$. Black lines indicate the blocks of the partition.
The 50% Cartesian credible set is shown in Figure 7. Number of legumes per plant is clearly very important for yield (posterior probability of inclusion is 0.96) and insertion of the first pod has some evidence (posterior probability of inclusion is 0.71). For all other variables, there’s no support for the inclusion of the variables.
7.3 Student performance
Cortez2008UsingDM describe a study of student performance. The response is the final marks of students in Maths or Portuguese in two Portuguese schools. There are 30 covariates covering family background, social habits and attitudes to education. The full list of regressors are included in Appendix D. There are a mix of numeric (continuous), ordinal, binary and categorical covariates. The two responses were modelled using a linear regression with categorical covariates included using dummy variables with the first category used as a baseline. As is common in data analysis, each regressor (including dummy variables) could be individually selected rather than allowing grouping of dummy variables related to different levels of the same variable.
Figure 8: Student performance example: $50\%$ Cartesian credible set with $M=2$ when final mark in Maths is used as the response. Black lines indicate the blocks of the partition.
The 50% Cartesian credible set when the response is the final marks in Maths is shown in Figure 8. Three variables have strong support as important: the number of failures in past classes (24, PIP = 1.00), whether the student receives extra educational support (25, PIP = 1.00) and the number of absences (39, PIP = 0.93). It seems reasonable that these variables should have an effect on educational performance. There is also some support for the inclusion of “Father’s job is other” (16, PIP = 0.75) and mother’s job is healthcare (9, PIP = 0.62). Two blocks of variables are also identified. Firstly, the difference between baseline (teacher) and several levels of mother’s job (civil service, or “at home”) (10, 11). The block PIP is 0.91. The highest probability submodel only includes “at home” (11, marginal posterior probability is 0.64). The second block is how often a student goes out with friends and their weekend alcohol consumption (35, 37). The posterior probability of including both is 0.92. It’s not surprising that these two variables are hard to disentangle since socialising can involve alcohol consumption. The highest probability submodel only includes going out (35, marginal posterior probability is 0.62).
Figure 9: Student performance example: $50\%$ Cartesian credible set with $M=2$ when final mark in Portuguese is used as the response. Black lines indicate the blocks of the partition.
The 50% Cartesian credible set when the response is the final marks in Portuguese is shown in Figure 9. Some variables which are important for Maths are also important for Portuguese: the number of failures in past classes (24, PIP: 1.00), whether the student receives extra educational support (25, PIP: 0.99), the number of absences (39, PIP: 1.00), and father’s job is “other” (16, PIP: 0.81). The block related social activities in Maths results expands to include sex (2) and the daily alcohol consumption (36) (the block PIP is 0.95). There’s a lot of uncertainty over the sub-model with largest posterior probability for a single sub-model being 0.22. In addition, the school (1, PIP: 1.00), age (3, PIP: 0.96), mother’s education (7, PIP: 0.95), whether the students wishes to take higher education (30, PIP: 1.00), and study time (23, PIP: 0.86). There’s also a block containing the reason for choosing the school (reputation, 17, or other, 19) which posterior probability of 0.72 that one of the variables is included.
@@ 7.4 GWAS example: Systemic Lupus Erythematosus
We consider a GWAS for Systemic Lupus Erythematosus using a case-control study. The presence or absence of the disease is the binary outcome in a logistic regression with the SNPs as explanatory variables. It is assumed that there are a number of causal SNPs which truly cause biological differences between the case and controls. A simple analysis would fit a logistic regression separately for each SNP and declare a SNP as associated with the outcome if it has a p-value smaller than a chosen (typically, extremely small) threshold. However, due to correlation in the genome, we usually see many SNPs declared as associated for each causal SNP. We refer to this group of SNPs as the signal. The goal of the analysis is to determine the number of independent signals and which SNPs are most likely the causal SNPs at each signal.
@@
The data contains 4,036 cases and 6,959 controls, with the cases and controls having European ancestry. In this example we only look at SNPs on Chromosome 3 ($42,430$ SNPs/variables). We apply Bayesian variable selection to the data using a logistic regression model with all SNPs as potential covariates and fit the model using an adaptive MCMC algorithm (Griffin and Wan, 2021). The full specification of the Bayesian model is provided in Appendix B. Figure 10 shows the PIPs for each SNP. There are several clear signals from SNPs with PIPs equal to 1 and others with large PIPs. The 50% Cartesian credible set is found by first retaining only those SNPs with a PIP above 0.04 (leaving 69 variables). The genomic distances of these 69 variables are given in Appendix D. The median model contains 13 variables which using the indices of the retained SNPs are 5, 10, 37, 45, 47, 48, 50, 51, 52, 53, 54, 58, 63.
@@
Figure 11: GWAS example: $50\%$ Cartesian credible set with $M=2$. Black lines indicate the blocks of the partition.
@@
The 50% Cartesian credible set is shown in Figures 11.. The most striking aspect is that the blocks of variables are contiguous which reflects the serial correlation of the SNPs. The results show that 11 variables in the median model do not include the possibility of exclusion in the credible set (10, 45, 47, 48, 50, 51, 52, 53, 54) and two variables are included as singletons with the possibility of exclusion but stronger support for inclusion (5 and 37). Two other variables from the median model are included as the most important submodels in blocks which have strong support for inclusion. These are 58 with 57 and 63 with 64, 65, 66, and 67. In addition, there is one block with strong support (and, which has no variables included in the median model) containing 59, 60, 61, and 62 (genomic distance: 159767453–159804903), which has block PIP equal to 0.84 (the posterior probability of modal sub-model for this block is 0.28).
@@ 8 Discussion
Cartesian credible sets are a useful way to summarize and visualize the output from Bayesian variable selection. The sets are constructed by finding a partition of the variable and an approximation to the posterior distribution where variables in different subsets are independent. Credible sets are defined as a Cartesian products of sets of sub-model and can be found using a simple heuristic algorithm which minimizes the KL divergence between the posterior and the approximation and chooses a parameter controlling the level of approximation and this can be chosen using a criterion which trades-off the accuracy of the approximation and overall the size of the credible set. The method can be run quickly as post-processing of an MCMC sample from the posterior distribution on the models. The results can be visualized by plotting the list of sub-models for each subset and coloured according to the posterior probability of the sub-models and the posterior probability of including at least one variable in a subset. This gives a much more complete picture of the posterior distribution than the usually plotted PIPs and helps understanding of the interchangeability of variables which are highly collinear. This allows us to avoid dilution effects from posterior mass being distributed across such variables by considering blocks of variables.
@@ A.1 Proposition 1
Appendix B Bayesian inference in regression models
B.1 Linear regression
@@
The posterior distribution was sampled using a simple Add-Delete-Swap sampler for $\gamma$ with Gibbs steps to update $\pi$ and $g$ and all other parameters integrated out. The samplers were run to collect 100 000 iterations with a burn-in of 100 000 and every 300th value recorded (total number of iterations 30 100 000).
B.2 Logistic regression
@@
The posterior was sampled using the ASI sampler for logistic regression from Griffin and Wan (2021). Three independent runs were used with each run containing 25 replicate chains. Each chain was run for 15 000 iterations with a burn-in of 5 000 iterations and the subsequent 10 000 samples thinned every 10th value. This gave 1 000 samples for each replicate chain, 25 000 samples for each run, and 75 000 samples overall.
Appendix C George and McCulloch (1997) example
Appendix D Further details of applications
The variable in the “behaviour of traffic in Sao Paulo, Brazil” example are shown in Table 4
No. Variable No. Variable No. Variable
27 Immobilized Bus 34 Incident Involving 41 Semaphore Off
Dangerous Freight
28 Broken Truck 35 Lack Of Electricity 42 Intermittent Semaphore
29 Vehicle Excess 36 Fire 43 15th December
30 Accident Victim 37 Point Of Flooding 44 16th December
31 Running Over 38 Manifestations 45 17th December
32 Fire Vehicles 39 Defect In The Network 46 18th December
of Trolleybuses
33 Occurrence Involving Freight 40 Tree On The Road
Table 4: The regressors for the Sao Paulo traffic example
In the GWAS example, the variable used after PIP screening are shown in Table 5
No. Distance No. Distance No. Distance No. Distance No. Distance No. Distance
@@ 33 Occurrence Involving Freight 40 Tree On The Road
The variable in the “student performance” example are shown in Table 6
Number Variable Type
1 school - student’s school (Gabriel Pereira or Mousinho da Silveira) Binary
2 sex - student’s sex (female or male) Binary
3 age - student’s age Numeric
4 address - student’s home address type (urban or rural) Binary
5 famsize - whether family size is less or equal to 3 Binary
6 Pstatus - parent’s cohabitation status (living together or apart) Binary
7 Medu - mother’s education Ordinal
(0 - none, 1 - primary education (4th grade), 2 – 5th to 9th grade,
3 – secondary education or 4 – higher education)
8 Fedu - father’s education Ordinal
(0 - none, 1 - primary education (4th grade), 2 – 5th to 9th grade,
3 – secondary education or 4 – higher education)
9–12 Mjob - mother’s job Categorical
(teacher, healthcare related, civil services, at home, other)
13–16 Fjob - father’s job Categorical
(teacher, healthcare related, civil services, at home, other)
17–19 reason - reason to choose this school Categorical
(close to home, school reputation, course preference, other)
20–21 guardian - student’s guardian (mother, father, other) Categorical
22 traveltime - home to school travel time Ordinal
(1 - $<15$ min., 2 - $15$ to $30$ min., 3 - $30$ min. to $1$ hour, or 4 -$>1$ hour)
23 studytime - weekly study time Ordinal
(1 - $<2$ hours, 2 - $2$ to $5$ hours, 3 - $5$ to $10$ hours, or 4 - $>10$ hours)
24 failures - number of past class failures (capped at 4) Numeric
25 schoolsup - extra educational support (yes/no) Binary
26 famsup - family educational support (yes/no) Binary
27 paid - extra paid classes within the course subject Binary
(Math or Portuguese) (yes/no)
28 activities - extra-curricular activities (yes/no) Binary
29 nursery - attended nursery school (yes/no) Binary
30 higher - wants to take higher education (yes/no) Binary
31 internet - Internet access at home (yes/no) Binary
32 romantic - with a romantic relationship (yes/no) Binary
33 famrel - quality of family relationships (1 - very bad to 5 - excellent) Ordinal
34 freetime - free time after school (1 - very low to 5 - very high) Ordinal
35 goout - going out with friends (1 - very low to 5 - very high) Ordinal
36 Dalc - workday alcohol consumption (1 - very low to 5 - very high) Ordinal
37 Walc - weekend alcohol consumption (from 1 - very low to 5 - very high) Ordinal
38 health - current health status (from 1 - very bad to 5 - very good) Ordinal
39 absences - number of school absences Numeric
Table 6: Variables used in the student performance example